OpenStudy (anonymous):

PLEASE HELP!!! WILL GIVE MEDALS!!! Racer A applies a 50 N force to a 2 kg race car. Racer B can only provide a force of 25 N. What should the mass of Racer B's race car be in order for both race cars to accelerate at the same rate? A. 4 kg B. 2 kg C. 1 kg

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@DanJS

OpenStudy (anonymous):

@ShadowLegendX

OpenStudy (anonymous):

@FAITH0_o

OpenStudy (shadowlegendx):

Racer A applies a 50 N force to a 2 kg race car. Racer B can only provide a force of 25 N. What should the mass of Racer B's race car be in order for both race cars to accelerate at the same rate? \[\frac{ 50N }{ 2kg } \rightarrow \frac{ 25N }{ ? }\]

OpenStudy (shadowlegendx):

\[\frac{ 25N }{ ? } \rightarrow \frac{ 50N }{ 2kg}\]

OpenStudy (shadowlegendx):

Force = mass x acceleration. a = F/m so since the acceleration is the same Fb/mb = FA/mA finally mb = mA x Fb/FA = 2kg x 25N / 50N =